If 39.9 g CuSO4 placed in a 250 mL, what will be the molarity of this solution?
1
Expert's answer
2017-11-07T15:57:07-0500
Molarity CM:
C_M=(n(CuSO_4))/(V(solution))
Where n(CuSO_4 )= (m(CuSO_4))/(M(CuSO_4))=(39.9 g)/(159.6 g/mol)=0.250 mol and V(solution)=0.250 L
C_M=(0.250 mol)/(0.250 L)=1.00 M
Answer: 1.00 M.